You are planning to deposit $100 in an account earning 6% interest rate.

a. How much money do you have at the end of three years if interest is compounded annually?

b. How much money do you have at the end of three years if interest is compounded semiannually?

c. How much money do you have at the end of three years if interest is compounded quarterly?

d. How much money do you have at the end of three years if interest is compounded monthly?

e. How much money do you have at the end of three years if interest is compounded daily?

USE FINANCIAL CALCULATOR AND SHOW WORK
Answer #1

FV = future value FOR FINANCIAL CALCULATOR

PV= present value USE THE TVM FUNCTION OF THE CALCULATOR

N= numbers of years HAVE TO CALCULATE THE FV

I/Y OR r = interest rate PV = -100 IN EVERY CASE

M = Number of compounding in a year

formula for FV

FV = PV ( 1 + r / M ) ^ NxM

a. How much money do you have at the end of three years if interest is compounded annually?

FV = ?

PV= 100   

N= 3

I/Y OR r = 6%

M = 1 (yearly)  

FV = (1.06)^3 x 100

= 119.10   

financial calculator steps use the TVM function

PV = - 100  

N= 3

I/Y = 6

press CPT AND THEN PV TO GET THE ANSWER

b. How much money do you have at the end of three years if interest is compounded semiannually?

FV = ?

PV= 100

N= 3

I/Y OR r = 6%

M = 2 (semiannually)

FV = (1 + 6 % / 2 )^(3x2) x 100  

= 119.405

financial calculator steps use the TVM function

PV = - 100  

N= 3 X 2 = 6 ( SEMIANNUAL COMPOUNDING)

I/Y = 6/2 = 3

press CPT AND THEN PV TO GET THE ANSWER

c. How much money do you have at the end of three years if interest is compounded quarterly?

FV = ?

PV= 100

N= 3

I/Y OR r = 6%

M = 4 ( quarterly)  

FV = (1 + 6 % / 4 )^(3x4) x 100

= 119.561

financial calculator steps use the TVM function

PV = - 100  

N= 3 X 4 = 12 ( quarterly COMPOUNDING )

I/Y = 6 / 4 = 1.5

press CPT AND THEN PV TO GET THE ANSWER

d. How much money do you have at the end of three years if interest is compounded monthly?

FV = ?

PV= 100

N= 3

I/Y OR r = 6%

M = 12 (monthly)  

FV = (1 + 6 % / 12 )^(3x12) x 100

= 119.668

financial calculator steps use the TVM function

PV = - 100  

N= 3 X 12 = 36 ( MONTHLY COMPOUNDING )

I/Y = 6 / 12 = 0.5

press CPT AND THEN PV TO GET THE ANSWER

e. How much money do you have at the end of three years if interest is compounded daily?

FV = ?

PV= 100

N= 3

I/Y OR r = 6%

M = 365 (daily)  

FV = (1 + 6 % / 365 )^(3x365) x 100

= 119.7199

financial calculator steps use the TVM function

PV = - 100  

N= 3 X 365 = 1095 ( DAILY COMPOUNDING )

I/Y = 6 / 365 = 0.016438  

press CPT AND THEN PV TO GET THE ANSWER